Write an example of a function whose derivative can be found by using the following rules:

a) Product rule and special function differentiation rules

b) Power rule, quotient rule, and chain rule

c) Chain rule twice

d) Implicit differentiation and special function differentiation rule



1
Expert's answer
2022-02-24T06:16:51-0500

Solution:

(a):

"f(x)=(2x+1)(3x-4)"

(b):

Power rule: "f(x)=x^7"

quotient rule: "f(x)=\dfrac{x+4}{4x-1}"

chain rule: "f(x)=\sin^4x"

(c):

Chain rule twice: "f(x)=\sin^4x^3"

(d):

"f(x)=y=(7x+2)^3(2x-5)^2"